Find f(x) and g(x) so that the function can be described as y = f(g(x)).
y = 4/(x^2) + 9

Answer by ikleyn(52835)

.
1.  g(x) = x^2,

    f(x) = (4/x) + 9.


or


2.  g(x) = 1/x,

    f(x) = 4x^2 + 9.


I was happy to see you followed my instructions on using parentheses.